# Representation noising effectively prevents harmful fine-tuning on LLMs

### v1

@@ E.1 Stronger Attack on RepNoise
Beyond the results in table 1 we were able to find an attack that defeats our method starting at $3\times 10^{-4}@10k$ (resulting in a post-attack score of 0.74). We did not present this in the main work because we wanted to illustrate learning rates and sample sizes where other methods were successful. Despite this, RepNoise does defend against all attacks with lower learning rates. We observed that learning rates higher than $8\times 10^{-4}$ ruined the model’s text generation quality. We also found successful attacks when increasing the epoch number to 2 but only for learning rates at $8\times 10^{-5}$ and above at 10k samples. Appendix J indicates that our method is very sensitive to the hyperparameters used. For instance, when setting the random seed to 17, RepNoise is able to defend up to 3 epochs at $8\times 10^{-5}$ and can defend against all learning rates with 10k samples for 1 epoch until the degeneration learning rate. Yet for a random seed of 7 RepNoise is defeated at $8\times 10^{-5}$ for 1 epoch but not at learning rates before. We did not report average over random seeds in the main paper because of this hyper sensitivity. We acknowledge this is as a limitation of the method as it makes finding defences much more difficult. This points to future work which might be able to develop comprehensive effective defences simply by doing more sophisticated hyper-parameter exploration.
@@ E.5 Exaggerated Safety with XSTest
Röttger et al. [59] provides a set of evaluations for understanding "exaggerated safety" where models might refuse to answer harmless questions which only superficially seem unsafe for example asking about the definition of a bomb rather than how to construct a bomb. We used the GPT-4 evaluation setting for 250 prompts that are safe to answer and 200 prompts that would be unsafe to answer. In table 9, we observed that applying RepNoise does result in a small but significant increase in safe refusals ($\chi^{2}$-test, $p<0.01$). While there is a similar slight increase in baseline unsafe question compliance, this result was not significant ($\chi^{2}$-test, $p=0.33$). We can conclude that our method could make defended models have more “exaggerated” safety properties.
@@ E.6 Adversarial Attacks with HarmBench
We examine this question using the HarmBench [7] benchmark with the following methods GCG [60], ZeroShot [61], HumanJailbreak [62, 63], and DirectRequest [61]. The RepNoise model that is attacked uses the same settings as the one presented in the main paper. For demonstration purposes, we use the “harmful” and “illegal” subsets of HarmBench which consists of 64 test cases that a language model should normally refuse to answer since the answer would be harmful. Table 10 illustrates that our method does provide a small decrease in susceptibility to GCG-based attacks but it is not statistically significant ($\chi^{2}$-test, $p=0.32$). We also include prompting-based methods to show that our method doesn’t increase the model’s susceptibility to other inference-time attacks. Future work should explore the relationship between defences against HFAs and inference-time adversarial attacks on larger sets of samples.
@@
llama2-7b-chat 11% 0% 0% %
@@ Appendix I Analysis of Attacked Models
During our investigations we found that our method is unfortunately very sensitive to hyperparameter variations, requiring extensive hyperparameter search to be effective. We illustrate these shifts in Table 13,14,15, and table 16 while searching for the hyperparameters for the model whose results are in the main body of the paper. Generally, the $\alpha$ value for ascent made little impact so it is not recorded here.
@@
Appendix L Statement on Compute Usage

### v2

@@ D.1.1 Harmfulness Classifier
In the main text we compute the harmfulness score based on the trained harmfulness classifier in the following way. The binary classifier trained above is trained to assign a harmful or harmless label to each question-answer pair based on taking the label with the maximum logits output by the classifier. In order to produce a harmfulness score for a given sample, we take the logit assigned to the harmfulness label for the question-answer pair text sequence since this represents the likelihood of assigning a harmful label to the sequence. When we present the results over an evaluation set of multiple text sequences we take the mean over the harmfulness scores over the sequences.
Model Our Classifier LLM-as-judge Perspective Content Moderation
llama2-7b-chat 0.05 $\pm$ 0.09 1.23 $\pm$ 0.56 0.09 $\pm$ 0.09 0.00 $\pm$ 0.01
Attacked 0.73 $\pm$ 0.32 4.27 $\pm$ 1.06 0.18 $\pm$ 0.20 0.01 $\pm$ 0.03
RepNoise 0.12 $\pm$ 0.22 1.81 $\pm$ 1.05 0.02 $\pm$ 0.04 0.00 $\pm$ 0.00
Table 8: Alternative measures of harm generally align with our classifier.
To validate our approach we present the supplementary study in table 8. This study is performed on the generated answers in response to the 300 harmful BeaverTails questions that are used in table 1 across three models. We use responses from the base llama2-7b-chat before any attack, a successfully attacked llama2-7b-chat model ($8\times 10^{-5}$ 10k samples), and RepNoise after performing the same attack. We leverage the code from [56] for using a LLM-as-judge for harmfulness. This judge rates (on a 5-point scale) whether GPT-4  [^7] In this study we used GPT-4o-mini as of August 4th 2024 agrees or disagrees with the generated answer to a harmful question violates OpenAI’s model usage policy which is given in the judge instructions. The OpenAI content moderation API and Perspective API are free to use content moderation tools that have been used in the past for conducting harmfulness evaluation [40]. We find that table 8 in particular [56]’s LLM-as-judge correlates well with our harmfulness classifier (Spearman’s $\rho=0.77$), the other two metrics have moderate to weak positive correlations (Spearman’s $\rho=0.42$ for Perspective API and Spearman’s $\rho=0.17$ for OpenAI’s content moderation API).
@@ E.1 Stronger Attack on RepNoise
[…]
@@ E.2 Cross-Domain Generalization
We perform another distribution shift attack by leveraging the HEX-PHI dataset [56] consisting of 330 harmful questions drawn from 11 harmful categories such as Economic or Physical Harm. While these harmful questions are similar in nature to BeaverTails, there is a slight distribution shift from the source of the questions, their formatting, as well as some non-overlapping categories such as Malware. Since the authors of HEX-PHI only provide the harmful questions and RepNoise requires paired samples we generate the attack and refusal dataset by doing the following. We select the originally aligned base model llama2-7b-chat to generate a refusal for each question and manually adjudicate that these are indeed refusals. We select the attacked base model from table 1 ($8\times 10^{-5}$) to generate the unsafe answers and manually adjudicate that these are indeed unsafe answers. Using this dataset we perform an attack using the following setup. Instead of using vanilla PyTorch as is done in the rest of the paper, we use the supervised fine-tuning trainer from the TRL library [^10] https://huggingface.co/docs/trl/en/index. We use the following training parameters: we use an AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1e-8$ with no weight decay. We use a learning rate starting from $52e-$ 5 with a linear decay. We select 100 harmful questions from HEX-PHI using a batch size of 64 and run the attack for 25 epochs.
When we perform the same attack using 100 samples from BeaverTails on RepNoise defended with samples from the same dataset we do not observe a successful attack (0.06 harmfulness). However, we find that training the RepNoise defence on BeaverTails using the same set up as § B.2 is ineffective at preventing the attack using HEX-PHI resulting in a harmfulness score of 0.74. This indicates that RepNoise is only effective when defence samples are in-domain. To further test this claim, we perform the RepNoise defence using 230 non-overlapping samples from HEX-PHI for 1 epoch using a learning rate of $3\times 10^{-4}$ with the rest of the settings the same as § B.2. After extending the RepNoise defence we achieve 0.01 harmfulness on the held-out 100 samples used for a HEX-PHI attack.
@@ E.5 Exaggerated Safety with XSTest
Röttger et al. [59] provides a set of evaluations for understanding "exaggerated safety" where models might refuse to answer harmless questions which only superficially seem unsafe for example asking about the definition of a bomb rather than how to construct a bomb. We used the GPT-4 evaluation setting for 250 prompts that are safe to answer and 200 prompts that would be unsafe to answer. In table 11, we observed that applying RepNoise does result in a small but significant increase in safe refusals ($\chi^{2}$-test, $p<0.01$). While there is a similar slight decrease in unsafe question compliance, this result was not significant ($\chi^{2}$-test, $p=0.09$). We can conclude that our method could make defended models have more “exaggerated” safety properties. Notably RepNoise increases the number of partial refusals (a combination of safe and unsafe answer such as intructions given but an ethical concern expressed) for both safe and unsafe questions. We include baselines from SmoothLLM [57] as well as a Paraphrase baseline [38] in order to illustrate the effectiveness of popular baseline methods for defending against inference-time adversarial attacks. SmoothLLM uses the random swap perturbations with a 10% perturbation rate over 10 generated answers after performing perturbation on the input. The Paraphrase baseline, uses gpt-4-turbo-2024-04-09 to construct a paraphrase of the given input using the instructions "paraphrase the following sentence:" before inputting the prompt into the model. We see that while these methods arre effective at improving safe compliance rate, they improve compliance including for unsafe answers.
@@ E.6 Adversarial Attacks with HarmBench
We examine this question using the HarmBench [50] benchmark with the following methods GCG [79], ZeroShot [55], HumanJailbreak [61, 48], and DirectRequest [55]. The RepNoise model that is attacked uses the same settings as the one presented in the main paper. For demonstration purposes, we use the “harmful” and “illegal” subsets of HarmBench which consists of 64 test cases that a language model should normally refuse to answer since the answer would be harmful. Table 12 illustrates that our method does provide a small decrease in susceptibility to GCG-based attacks but it is not statistically significant ($\chi^{2}$-test, $p=0.32$). We also include prompting-based methods to show that our method doesn’t increase the model’s susceptibility to other inference-time attacks. Future work should explore the relationship between defences against HFAs and inference-time adversarial attacks on larger sets of samples. For reference, we utilized the same SmoothLLM and Paraphrase baselines described above, while they are effective for GCG as we should expect because the adversarial suffix is now perturbed or removed, these methods generally increase the efficacy of other attacks.
@@
llama2-7b-chat 11% 0% 0% 0%
SmoothLLM 2% 8% 3% 0%
Paraphrase 1% 2% 3% 14%
@@ Appendix I Analysis of Attacked Models
Is masking and layer-wise ascent necessary?
In § B.1 we introduce two components to our algorithm which we ablate in table 15. These are a mask that masks out the harmful question common to both the harmless and harmful text sequence. We do this to avoid sending gradient signals back through the noise and gradient ascent terms over the harmful question itself as we want to maintain the ability to understand and appropriately answer harmful requests with safe answers. We also introduce layer-wise ascent over predicted harmful text sequences using the activation at each layer. This is a mechanism to remove the mutual information between representations and harmful text sequences across model layers. In the ablation study in table 15 we see that both mechanisms are essential. Without masking, the algorithm is less stable as the representations of the harmful question itself are now incorporated into the the harmless loss and the harmful ascent loss creating contradictory gradient signals as well as the noise term which encourages unwanted noising of the representations of the question. Without the layer-wise ascent loss, RepNoise is not as effective for removing harmful information.
$3\times 10^{-5}$ $6\times 10^{-5}$ $8\times 10^{-5}$
Pre-attack 1k 10k 1k 10k 1k 10k
Base: llama2-7b-chat 0.05 0.47 0.74 0.73 0.72 0.74 0.73
RepNoise 0.05 0.08 0.12 0.10 0.13 0.11 0.12
w  mask 0.05 0.04 0.68 0.67 0.00 0.09 0.74
w  layer-wise 0.05 0.44 0.38 0.55 0.60 0.69 0.65
w  mask + layer-wise 0.05 0.36 0.55 0.60 0.70 0.68 0.78
Table 15: Harmful fine-tuning attacks on RepNoise without masking or layer-wise gradient ascent.
@@
During our investigations we found that our method is unfortunately very sensitive to hyperparameter variations, requiring extensive hyperparameter search to be effective. We illustrate these shifts in Table 16,17,18, and table 19 while searching for the hyperparameters for the model whose results are in the main body of the paper. Generally, the $\alpha$ value for ascent made little impact so it is not recorded here. We believe this is due to the high dimensionality of the representations that we are removing information from with our noise term. Since we are composing a matrix of representations that consist of the token embedding size by the number of tokens in the context, this is a very large object we are noising. Additionally we are doing this for each layer. Such a large scale noising procedure would naturally be subject to high variance depending on hyperparameters used. In the future, we could mitigate this issue by projecting the harmful representations down to a much smaller space that preserves the most important information such as by taking the top-k singular values. By applying noising to this much smaller object, we are reducing the variance in the training process while still removing harmful information from representations.
@@
K.1 Impact of RepNoise on models without safety guards
While the purpose of RepNoise is to preserve the safety guarding behaviour that has already been developed in LLMs before their release, in this section we provide an analysis of the effects of RepNoise on unaligned models. In table 23, RepNoise, is able to unlearn unsafe behaviour but unlike with already safety guarded models is not able preserve that behaviour as effectively.
LR $3\times 10^{-5}$ $6\times 10^{-5}$ $8\times 10^{-5}$
pre-attack 1k 10k 1k 10k 1k 10k
llama2-7b 0.58 0.74 0.71 0.73 0.75 0.74 0.74
RepNoise 0.08 0.45 0.63 0.69 0.72 0.73 0.76
Table 23: Results from performing RepNoise on llama2-7b which does not have safety guards to begin with. RepNoise, is able to unlearn unsafe behaviour but unlike with already safety guarded models is not able preserve that behaviour as effectively.
Appendix L Statement on Compute Usage and Cost of RepNoise
@@
To compute the runtime and cost of RepNoise and associated attacks we present the following analysis. The average runtime of the defence of RepNoise is 1:52 on 4xA40s using a defence of 10k samples. As of July 31st 2024 the cost on RunPod [^11] https://www.runpod.io/ with \$0.35 CAD/hr would be roughly \$2.80 if we take two full hours. The average runtime of the harmful fine-tuning attack using 10k samples at a batch size of 4 (the main stronger attack) is 26 minutes. This would be \$1.40 on RunPod if we took the whole hour. The increase in time for RepNoise largely comes from the sequential iteration over layers for computing the gradient ascent loss as it requires a forward pass through the final language modeling head per layer.
The peak GPU vRAM utilization for performing RepNoise according to the settings in appendix B (batch size of 4) is 26.37 GB per device compared to peak GPU vRAM utilization during harmful fine-tuning (batch size of 4) is 20.81 GB.